OpenStudy (anonymous):

How can I turn this into a fraction... (1-3x)^(3/x)

OpenStudy (anonymous):

take the log first

OpenStudy (anonymous):

would that become \[(\frac{ 3 }{ x })\ln(1-3x)\]

OpenStudy (anonymous):

or else use the definition \[(1-3x)^{\frac{3}{x}}=\exp\left(\frac{3\log(1-3x)}{x}\right)\] the work will be identical in either case

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

then use l'hopital , then whatever you get for the limit, raise e to that power

OpenStudy (anonymous):

Ahh I see thanks so much! :D

OpenStudy (anonymous):

actually now that i look, it doesn't say that this is a limit question, but i am almost certain it is right?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Haha yep! It is :D one of the l'hopital rule limits xP

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

Smarty pants :P

OpenStudy (anonymous):

lol